OpenStudy (anonymous):

How would I simplify 1-2sinx=-cosx to the point where I could solve for x on [0,2pi)

OpenStudy (amistre64):

1 = sin^2+cos^2

OpenStudy (amistre64):

1-2sinx=-cosx sin^2+cos^2 -2sin = -cos (sin^2-2sin) = - (cos^2 +cos)

OpenStudy (anonymous):

I got sinx(sinx-2)=-cosx(cosx+1)

OpenStudy (anonymous):

I don't see where I could go from here except for maybe getting a -tanx.

OpenStudy (amistre64):

one idea is to complete the squares sin^2-2sin = - (cos^2 +cos) sin^2-2sin +1 -1 = - (cos^2 +cos +1/4) + 1/4 (sin -1)^2 -1 = - (cos + 1/2)^2 + 1/4 (sin -1)^2 + (cos + 1/2)^2 = 5/4 \(\frac{4}{5}\)[(sin -1)^2 + (cos + 1/2)^2] = 1 jsut a thought tho :)

OpenStudy (amistre64):

think of a circle centered at (1,-1/2) with a radius of sqrt(5)/2

OpenStudy (anonymous):

?

OpenStudy (anonymous):

Can you put hat into the equation editor thingy please?

OpenStudy (amistre64):

\[(\sin x -1)^2 + (\cos x + 1/2)^2 = \frac54\]

OpenStudy (anonymous):

Wait, but how does that get me closer to solving for x?

OpenStudy (amistre64):

let u = sin(x), let v = cos(x) but maybe you want a different method

OpenStudy (anonymous):

:? I'm so lost now.

OpenStudy (anonymous):

@cwrw238